OpenStudy (anonymous):

PLEASE HELP Use implicit differentiation to find an equation of the tangent line to the ellipse x^2/2 + y^2/72 = 1 at (1,6)

OpenStudy (anonymous):

is someone actually going to reply...?

OpenStudy (kc_kennylau):

\[ \frac d{dx}\frac{x^2}2+\frac d{dx}\frac{y^2}{72}=\frac d{dx}1\\ \frac{x^3}6+\frac d{dy}\frac{y^2}{72}\times\frac{dy}{dx}=0\\ \frac{y^3}{216}\times\frac{dy}{dx}=-\frac{x^3}6\\ \frac{dy}{dx}=-\frac{x^3}6\times\frac{216}{y^3}=-\frac{36x^2}{y^2}=-\frac{36x^2}{y^2} \]

OpenStudy (kc_kennylau):

@lamp12345 sorry I saw @G-unit was replying so I didn't reply

OpenStudy (anonymous):

@kc_kennylau thats not an answer choice: -12x+8 -5x+10 -10x+8 -8x+10 -6x+12

OpenStudy (kc_kennylau):

Well it's not finished

OpenStudy (kc_kennylau):

you have to substitute x=1 and y=6 to find the slope of the tangent

OpenStudy (anonymous):

oh ok

OpenStudy (kc_kennylau):

1. Substitute x=1 and y=6 to the derivative to find out the slope 2. Substitute x=1 and y=6 to the original function to find out the point cut by the tangent. 3. Point-slope form 4. Convert it to slope-intercept form by making y the subject

OpenStudy (anonymous):

ok i did that and didn't get an exact answer but it was closest to -6x+12.

OpenStudy (anonymous):

Okay, first we want to take the derivative of both sides. d/dx (x^2)/2 + (y^2)/72 = d/dx 1 d/dx (x^2)/2 + (y^2)/72 = 0 So the derivative of x^2/2 is x. If you need that explained as well let me know but it's pretty basic. x + d/dx (y^2)/72 = 0 Here's where it might get a little confusing. Let's go ahead and isolate the d/dx (y^2)/72 so it doesn't get messy. When you take the d/dx of a y variable, here's the rule. You basically treat it like an x, and multiple it by dy/dx So it becomes dy/dx(2y) * 1/72 Or simplified to (1/36y) dy/dx Now we plug it back in to what we already had: x + (1/36y) dy/dx = 0 Move the x to the other side (1/36y) dy/dx = -x Divide by 1/36y dy/dx = -36x/y Substitute in the x=1 and y=6 -36(1)/6 =-6 is the slope of the tangent line Now do the tangent work, and you should get the correct answer. I think Kenny's answer was off which is why you didn't get the correct result.

OpenStudy (anonymous):

@G-unit so my answer was correct? -6x+12

OpenStudy (anonymous):

oh wait you said I didnt get the correct result. whats the tangent work I have to do @G-unit

OpenStudy (anonymous):

y - 6 = -6 (x-1) y - 6 = -6x + 6 y = -6x + 12 Yes, it is the answer.

OpenStudy (anonymous):

thanks!